# PSAppendix - Selected Primary Studies

### v1

@@ 4.3.9. Domain experts
These are the people who have expertise and knowledge about the problem domain, e.g., clinical neurologists for an ML-enabled psychological monitoring system [174], urban planners for a smart city system [30]. Domain experts play a crucial role in domain discovery [41], engineering of domain specific requirements [24], and fulfilling business goals [41], [117] due to their deep understanding of the problem domain and environment. They are “irreplaceable actor within ML projects” [24]. Moreover, they also take part in collecting, cleaning, and labeling data, interpreting data distributions, developing and validating hypotheses about data, model evaluation and interpretation of the results [32], [151], [111]. According to the literature, domain experts either collect and process data and transfer it to the MLOps team, or they are directly accessible to the MLOps team for collaboration, especially with data engineers, data scientists and ML engineers [144], [174], [173].
4.3.10. Business analyst and experts
Analyzing the problem that needs to be solved, defining clear objectives, and assessing whether an ML-enabled system could answer the problem are the primary tasks of these people [86], [175], [154], [93], [97]. The MLOps team collaborates with business experts to gain a better understanding of the problem and convert it to a successful ML solution that aligns with the business goals. Depending on the domain and project organization, business experts’ and domain experts’ responsibilities may refer to the same tasks; both help people to understand the problem (business) domain.
4.3.11. Collaborative activities among roles
@@
In a typical MLOps pipeline (Figure 3), various roles collaborate with each other for various responsibilities. Data engineers, data scientists, and ML engineers are mainly responsible for carrying out data collection and preparation, data analysis, and model development and deployment stages, respectively. However, these roles often collaborate with business and domain experts. Particularly, data engineers work together with data scientists and ML engineers to decide on the appropriate data format necessary for further steps. Data scientists visualize the data analysis results and receive feedback from ML engineers on the potential model performances. Data scientists also provide insights to ML engineers regarding model design and optimization. Data engineers, scientists, and ML engineers often collaborate with business and domain experts to interpret the data and model. ML engineers and sometimes data scientists accompany the MLOps engineers at the initial monitoring to ensure the stability of the model and step in when the model or features need optimization. On the other hand, MLOps engineers interact with every role to ensure the optimization of the automated pipeline, the operational stability of the ML system, and the smooth collaboration across various roles.
@@ 5.1. Adapting Team Structure and Work Practices for MLOps
Integrate practices for cross-silo concerns across ML life cycle: Cross-silo concerns include AI fairness, AI explainability, and security. Instead of addressing these concerns at any particular phase, they should be a cross-silo effort, which each part of the life cycle implementing different aspects. Amazon AI Fairness and Explainability Whitepaper [152] recommends integrating bias measurement before-, during-, and after-training, as well as in production. The detected bias then can be addressed using pre-processing, training tuning, and post-processing.
Security is another cross-silo concern of ML life cycle. Zhang and Jaskolka [105] introduce the concept of Secure Machine Learning Operations (SecMLOps) based on the People, Processes, Technology, Governance, and Compliance (PPTGC) framework. SecMLOps embeds security controls across the whole life cycle. In the requirement phase, security objectives and requirements of ML models and data are defined, and security models are established. These models drive the development of security policies regarding data, model, physical security, personnel security, administrative security, and network security across the ML life cycle and all involved parties. These models also drive the development and integration of security monitoring and incident response components that monitor ML models in production.
@@ 5.5. Systemizing Model Deployment
Using software container to package and deploy ML code:
A common practice is packaging ML code with all of its dependencies within software containers, and then applying optimizations such as parallelization and load balancing upon the containerized workloads [179], [172]. Containerized workloads can also be operated at scale by software container orchestrate platforms such as Kubernetes [156]. By encapsulating ML models as docker containers practitioners can reap the benefits of container management platforms such as AWS Fargate which provides serverless options for ML model deployment [111]. Kubernetes enables automated deployments, monitoring, and scaling of ML models to ensure efficient operation of deployed ML models, thus overcoming limitations of manual model deployments [128]. Syrigos et al. [128] propose an extension for Kubernetes Scheduler where container scheduling decisions are made in a latency-sensitive manner in addition to CPU and RAM of the computing resources to deploy ML models.
@@ 5.6. Setting up Automated Pipelines for Cross-silo Activities
Developing and utilizing automated pipelines: MLOps pipelines can be implemented with various tools. At a basic level, software scripts (e.g., Bash or Python) for submitting training or testing jobs can form a rudimentary version of automated pipelines. Moving up a level, organizations can adapt the existing tools and platform for CI/CD to MLOps workflows. GitHub Actions, CircleCI, ArgoCD, and Jenkins were among the tools suggested by the existing literature and industry best practices for adaptation to MLOps. For instance, Stieler and Bauer [136] proposed a Git workflow for an “active learning life cycle”, which contains four interconnected iterations of data management, model training, software development, and system operations. In the proposed workflow, a CI/CD tool hosted on a resource-rich machine with GPUs executes ML pipelines for training, testing, and integrating ML models based on the code committed by developers to a source control management system. Alternative, generic workflow engines such as Apache Airflow [^6] https://airflow.apache.org/use-cases/mlops/ can be adopted to MLOps pipelines. For instance, the FLScalize [172], an MLOps platform for federated learning, utilizes Airflow to orchestrate the interaction sequences among federated learning servers and clients as well as other monitoring workflows. Finally, there have been more ML-specific workflow orchestration tools (e.g., SageMaker Pipelines, IBM Watson Pipelines, Google Vertex AI Pipelines, Netflix Metaflow, ZenML) for defining and running MLOps pipelines.
@@ 5.7. Abstracting MLOps Complexity
6. RQ3: What are the Challenges Practitioners Face when Adopting MLOps?
@@
In RQ2 we discuss work practices, techniques and technologies that organizations need to implement to enable MLOps. However, there are challenges that hinder teams from efficiently adopting MLOps into their organization’s ML-based software system productionalization process. Thus, we identify such challenges and highlight reported solutions from the literature, and summarize them in Table 5.
Table 5. MLOps adoption challenges and proposed solutions
MLOps Adoption Challenges Primary Aspects Proposed Solutions
Socio-technical Challenges Cultural Changes: Disruptions caused by MLOps related culture and work practices in traditional software management process. Openness and willingness of teams to change [103] Continuous collaboration among team members from multiple disciplines [104], [108] Available MLOps processes, practices and tools to promote collaboration [104], [108] [103], and identify/resolve collaboration conflicts [104]
Successful implementation of cultural changes [103] Follow a human-centric approach to implement cultural changes [154]
MLOps Fragmentation: Fragmentation of practices, technologies and tools due to rapid growth in MLOps. Inherent complexity of MLOps [154], [132] MLOps Standardization [154]
Accidental complexity due to fast evolution of MLOps [154] Creating domain-specific MLOps reference architectures [134], [133], [127]
Responsible MLOps: Challenges related to achieving regulatory compliance, AI fairness, transparency, accountability and security throughout the MLOps life cycle. Lack of expertise and standardized MLOps practices related to compliance, fairness, explainability and security [105], [121] Frameworks integrating responsible AI practices into MLOps [108], [117]
Limitations in end-to-end MLOps tools for facilitating responsible AI [152] [114] Tools for enabling and automating responsible AI practices across ML lifecycle [152], and their detailed documentation [105]
MLOps Pipeline Related Challenges Complexity : Complexity of ML pipelines hinders their use during initial stages of the ML development High setup and maintenance overhead. [129] Frameworks for creating lightweight ML pipelines. [129]
High level of expertise required to operate pipelines. [129]
High infrastructure costs [129]
Optimized deployment : Optimal placement of pipeline modules within distributed environments. Operation across distributed and heterogeneous resources [127], [128] Strategic modularization [127]
Manual provisioning of pipelines in distributed and dynamic environments is error prone and sub-optimal [127; 128] Dynamic provisioning of containerized pipeline modules [127]
Training challenges due to ML models growth [126] Dynamic resource provisioning for parallel or distributed model training [126]
Managing large scale pipelines : Maintaining thousands of production pipelines without failures. Failure-prone, recurring and inter-dependent production pipelines. [131] [176] Automation of pipeline operation monitoring and failure root cause analysis. [131] [177]
Manual, time-consuming monitoring and resolution of failures. [131] [177]
Continual learning : Operation of efficient continual learning (online learning) pipelines for high velocity and high volume streaming data Synchronization between learning, serving and maintenance activities [124], [125] Horizontal and vertical scaling of online-learning tasks based on predictions demand, data velocity, and volume [125]
Overwhelming model updating due to incoming high-velocity data streams [125]
Automation Decision Dilemma : Deciding which activities of the pipelines to automate and which needs human involvement Automation of certain ML life cycle activities (model development, model validation) reduces incorporation of human judgment which can negatively effect the use of responsible AI [110], [114] Design and implementation of pipelines in which the tedious tasks are automated and human insights are incorporated wherever it is needed throughout the lifecycle [110], [114] [117]
MLOps Platform Related Challenges Platform Customization : Decision on whether to use a managed MLOps platform or create a customized platform. Complexity and cost of managing own platforms add a heavy burden [108] Building managed platforms with integration and customization support [108]
Cloud-based managed platforms lacks fine-grain control and customization support [108] Adopt a modular architecture [108]
Increase support for open-source ML tools and their integration [108]
Scalability : Ability of the MLOps platforms to maintain smooth execution of MLOps services as the number of service requests increases. Lack of scalability as the number of service requests fluctuates in monolithic architecture [120] Using Microservice Architecture for platform design [120]
Artifact Management : Management of data, models and associated metainformation to ensure ML experiment tracking/governance and reproducibility Challenging versioning, traceability and reproducibility due to increased number of artifacts [85], [164] Tools/ platforms supporting version controlled and conjugated management of artifacts [122]
Reproducibility based on conjugated management of the related artifacts. [122]
Distributed computing environments : Challenges with design, deployment and maintenance of MLOps pipelines in decentralized edge environments Edge-AI has given rise to paradigms like TinyML, Federated ML at edge, which have MLOps related requirements different from the cloud-based platforms. [123] Build platforms that support MLOps support for usecases such as TinyML and Federated ML at the edge. [123]
Migration to MLOps platforms : Challenges faced by practitioners when integrating existing projects to MLOps platforms Tedious process of retrieving information from existing project that are required as input for the platform [178] Document the ML components and artifacts and use it for migration process [178]
Selecting a suitable MLOps platform for the project [178] Selecting a platform based on the architecture of the existing project [178]
Incompatibility between MLOps and DevOps tools [160] [19]
Use tools that offer hybrid CI/CD environment [178]
@@ 6.3. MLOps Platform Related Challenges
MLOps platforms are designed for end-to-end ML lifecycle management, as a means of abstracting the complexity of implementing MLOps as discussed in 5.7. This section highlights challenges and limitations related to creating and using such platforms covering platform provider and practitioner perspectives.
@@ 6.3.1. Platform customization
For companies looking to build and deploy ML models, the main goal is to generate business value from data. Thus, for most of the practitioners, the underlying MLOps platform is only instrumental and their aim is to minimize the effort on implementing and managing the platform. Cloud-based managed MLOps platforms like Sagemaker are advantageous to meet this requirement. However, using fully managed MLOps platforms means the practitioners would lose fine-grained control over the technology stack and the services supported by the platform may not meet specialized requirements of many businesses [108]. This drives practitioners to implement and manage their own platforms which add additional complexity and cost. Practitioner expectation in this scenario is easy customization of managed platforms by integrating open-source, proprietary or homegrown tools. To meet this requirement managed MLOps platform providers need to design and implement their platforms following a modular architecture, integration and customization support to enable practitioners to integrate open-source tools. As an example, AWS highlights the design of Amazon SageMaker which provides practitioners the flexibility to integrate their own workflows with open-source APIs supporting TensorFlow, PyTorch, Kubeflow, etc [108].

### v2

@@ 3.2.9. Team management
Domain experts possess deep knowledge of the problem domain, such as clinical neurologists for an ML-enabled psychological monitoring system [174] or urban planners for a smart city system [30]. They are described as “irreplaceable actors within ML projects” [24] due to their crucial role in domain discovery [41], engineering domain specific requirements [24], and fulfilling business goals [41], [117]. Moreover, they contribute to data collection, cleaning, labeling, hypothesis development, model evaluation, and result interpretation [32], [151], [111]. They either provide domain knowledge and processed data to the MLOps team in the early stages of a project and/or collaborate directly with data engineers, data scientists and ML engineers during the ML phases to validate data and model [24], [144], [174], [173].
Business analysts define the problem, set clear objectives, and assess whether an ML-enabled system can provide a solution to the problem at hand [86], [175], [97]. MLOps team collaborates with business analysts to gain a better understanding of the problem and convert it to a successful ML solution that aligns with the business goals. Their role may overlap with domain experts, as both contribute to understanding the problem domain. Business analysts also define business requirements, key performance indicators (KPIs), and model quality metrics, leveraging query processing on structured datasets to analyze business problems [86], [154].
Data engineers are mainly responsible for fundamental data activities, enabling data operations in a high-quality, secure, and scalable way [1], [85]. Data engineers’ typical set of tasks include identifying suitable data sources, verifying data legality, collecting data from various sources, cleaning them, and preparing labeled data [5], [51], [93]. This data engineering process is characterized as ”extract-transfer-load” which involves transferring data into consumable formats for data scientists and ML engineers [112], [106]. Beyond data preparation, data engineers set up and manage data storage infrastructures, build automated data pipelines [1], and enable big data storage solutions such as centralized data lakes [69] and streaming platforms [125]. Their role extends to security, privacy, quality, and governance of data [86], [106]. Some organizations define specialized roles, such as data curator, responsible for custom data acquisition [19], and data expert, applying noise reduction techniques [5].
Data scientists are mainly responsible for exploring and analyzing the data made available by data engineers, training, tuning, and evaluating models to derive insights [122], [85]. Particularly, their initial task is to understand the business problem and its domain, evaluate the feasibility of an ML-enabled solution [24], [114]. At this step, data scientists work with domain experts and business analysts to discuss objectives and key results (OKRs) [86]. Data scientists also conduct feature engineering to identify the best combinations of features for training effective ML models [104]. Tasks such as preparing training, test, and validation sets, choosing suitable ML algorithms and training approaches, and selecting appropriate evaluation metrics (e.g., accuracy) are also carried out by data scientists [112]. Moreover, as data engineers build the initial part of the ML pipeline, data scientists are responsible for designing and building scalable and high quality data analytics pipelines for experimentation and model training [114]. Additionally, they ensure the trackability of multiple versions of experiments, datasets, and models [122]. Also, they monitor the early results after the deployment, collaborating with business analysts and ML engineers, whether the model meets technical, business, and regulatory requirements [114]. If necessary, data scientists iterate on the model to improve its performance [93].
ML engineer and data scientist roles often overlap, with some studies suggesting they are interchangeable [135], [141]. However, many studies distinguish the role of the ML engineers with their primary tasks being model optimization, deployment, and monitoring [114], [154]. Whereas, experimenting with datasets and models, and optimizing the model and its parameters are tasks shared with data scientists [118], [26]. In particular, ML engineers contribute to monitoring and production activities, collaborating with data scientists and business stakeholders to ensure the deployed model meets performance and business requirements [114], [107]. They are also responsible for executing appropriate deployments, such as online serving, and monitoring the model’s performance in the production environment [114], [107].
MLOps engineers are responsible for building automated pipelines, maintaining operational stability of the entire ML system, and facilitating collaboration across roles [106], [117]. While data engineers, data scientists, and ML engineers design and implement the pipelines for data collection, analysis, model development, and deployment, MLOps engineers focus on building, optimizing, and integrating these pipelines, including services like cloud and containers [106]. They manage dependencies, versioning, tracking mechanisms, and platform governance. MLOps engineers also monitor data, models, and applications along with ML engineers [117], and step in when human intervention is needed—such as when re-training is required due to data shifts or resolving dependencies between data and models, [137]. Effective communication and collaboration management across roles during these tasks is also the responsibility of MLOps engineers [105].
DevOps engineers also play a role in ensuring the seamless deployment, integration, and stability of machine learning models within production systems. Their responsibilities often overlap with those of MLOps engineers, operational engineers, and software engineers [55], [69]. While some studies indicate DevOps engineers in traditional software projects are replaced with MLOps engineers for ML projects [109], a few mention DevOps and MLOps engineers work within the same team. For example, Neghawi et al. [117] report that while software engineers develop and validate applications, DevOps engineers construct, test, and deploy working systems, and MLOps engineers continuously monitor the entire system to ensure operational stability.
Software engineers develop the final applications and services that support ML production. Application developer is another common reference to the same role [1]. They often collaborate with data scientists and ML engineers in data processing and analysis, and model building [115], [151]. Besides, deployment, packing, releasing, configuring, and monitoring of ML systems are also reported as the role of software engineers [161], [144], [154]. Additionally, quality assurance engineers are responsible for testing the software application and services to ensure product quality [36].
Operational engineers or operational team are not consistently referenced across all studies. But some highlight their role as taking over after data scientists and/or ML engineers hand over the ML model [24]. They are responsible for deployment operations including packaging, releasing, and configuring ML systems [144] and continuous monitoring of the model’s performance [104]. They serve as a bridge between data science and software engineering teams, facilitating iterative improvements by reporting issues, providing feedback, and ensuring the model remains efficient and reliable in real-world applications [67]. In the absence of explicit mention of operational engineers, these tasks are carried out by ML [114], and sometimes software [144], and MLOps engineers [106].
Technical and solution architects are responsible for designing and building ML systems and services that serve as the backbone of ML operations in production [70]. Their role includes secure design, selecting appropriate technologies and design paradigms, tracking and monitoring models in production, and integrating cloud-based applications and web services [105], [155]. Also, referred as IT team or IT architects in the literature [105].
Security experts ensure the integrity, confidentiality, and availability of ML assets by defining security objectives and requirements at the project’s initiation [105]. Some studies refer to them as a dedicated security team [108] or security engineers [29], while others define them as data and software engineers with security expertise [105]. Their role involves securing the overall ML system, addressing potential attacks during data ingestion and feature engineering, and enforcing secure coding practices [105].
Compliance teams focus on mitigating unfairness and bias in ML systems throughout the model training, deployment, and monitoring phases [108], [152]. Recently, risk professionals have become involved in MLOps projects to manage data and model-related risks [155]. Specifically, model risk manager is responsible for ML model inventory management, model explainability, and pipeline management [111], [106], while other studies only mention risk professionals for managing any data and technology related risks [155]. Data steward and data quality board specifically focus on data quality management [129], and data governance officer focus on data governance and privacy [114].
Business owners, business managers, and business stakeholders are not clearly distinguished in the literature yet. However, they are often associated with organization level decision making [85], supervision of other roles in development and operations [105], and evaluating the production of ML systems from an operational and regulatory perspective [69]. Product experts, owners, and managers are associated with defining and optimizing product features to enhance performance and usability. For example, product experts identify feature set candidates based on specific performance issues and apply feature engineering techniques alongside to improve model outcomes [31]. Product managers make data-driven decisions by reviewing relevant collected and visualized information [34].
@@
Collaborative activities among roles: In a typical MLOps pipeline (Figure 3), various roles collaborate to fulfill different responsibilities. Often, one role is the primary owner of an activity, while other roles provide support as collaborators [86]. Data engineers, data scientists, and ML engineers often collaborate with domain experts and business analysts during data collection and preparation, data analysis, and model development and deployment stages [106]. Particularly, data engineers work together with data scientists and ML engineers to determine the appropriate data format for subsequent steps [111]. Additionally, data engineers, scientists, and ML engineers often collaborate with business and domain experts to interpret both the data and the model [111]. ML engineers and occasionally data scientists accompany the MLOps engineers at the initial monitoring to ensure the model’s stability and step in when the model or features need optimization [117]. Meanwhile, MLOps engineers interact with every role to ensure the optimization of the automated pipeline, the operational stability of the ML system, and the smooth collaboration across various roles [105],[87].
MLOps is perceived as a multidisciplinary field integrating continuous engineering practices (CI/CD), DevOps principles, data science, machine learning, and software engineering to systematically manage the lifecycle, development, deployment, and operation of ML-enabled software systems. It encompasses traditional ML activities with core MLOps tasks such as continuous deployment and monitoring, while placing increasing emphasis on quality assurance and compliance and efficient management of pipelines and artifacts, and cross-role collaborations. Emerging roles such as MLOps engineers, data governance officers, and compliance teams illustrate increasing specialization within MLOps, emphasizing its multidisciplinary character. Findings of RQ1 reflect the evolution of MLOps into a more structured, mature, and systematically governed approach for operationalizing ML systems.
@@ 4. RQ2: What are the state-of-the-art practices and techniques proposed for implementing MLOps?
Figure 4. MLOps solutions and the corresponding ML productionalization challenges
@@ 4.1. Adapting Team Structure and Work Practices for MLOps
Integrate practices for cross-silo concerns across ML life cycle: Cross-silo concerns include AI fairness, AI explainability, and security. Instead of addressing these concerns at any particular phase, they should be a cross-silo effort, which each part of the life cycle implementing different aspects. Amazon AI Fairness and Explainability Whitepaper [152] recommends integrating bias measurement before-, during-, and after-training, as well as in production. The detected bias then can be addressed using pre-processing, training tuning, and post-processing. Security is another cross-silo concern of ML life cycle. Zhang and Jaskolka [105] introduce the concept of Secure Machine Learning Operations (SecMLOps) based on the People, Processes, Technology, Governance, and Compliance (PPTGC) framework. SecMLOps embeds security controls across the whole life cycle. In the requirement phase, security objectives and requirements of ML models and data are defined, and security models are established. These models drive the development of security policies regarding data, model, physical security, personnel security, administrative security, and network security across the ML life cycle and all involved parties. These models also drive the development and integration of security monitoring and incident response components that monitor ML models in production.
@@ 4.5. Systemizing Model Deployment
Using software container to package and deploy ML code: A common practice is packaging ML code with all of its dependencies within software containers, and then applying optimizations such as parallelization and load balancing upon the containerized workloads [179], [172]. Containerized workloads can also be operated at scale by software container orchestrate platforms such as Kubernetes [156]. By encapsulating ML models as docker containers practitioners can reap the benefits of container management platforms such as AWS Fargate which provides serverless options for ML model deployment [111]. Kubernetes enables automated deployments, monitoring, and scaling of ML models to ensure efficient operation of deployed ML models, thus overcoming limitations of manual model deployments [128]. Syrigos et al. [128] propose an extension for Kubernetes Scheduler where container scheduling decisions are made in a latency-sensitive manner in addition to CPU and RAM of the computing resources to deploy ML models.
@@ 4.6. Setting up Automated Pipelines for Cross-silo Activities
Developing and utilizing automated pipelines: might be able to save time by having their well-defined model validation pipelines approved, saving time on having the models themselves validated and approved. MLOps pipelines can be implemented with various tools. At a basic level, software scripts (e.g., Bash or Python) for submitting training or testing jobs can form a rudimentary version of automated pipelines. Moving up a level, organizations can adapt the existing tools and platform for CI/CD to MLOps workflows. GitHub Actions, CircleCI, ArgoCD, and Jenkins were among the tools suggested by the existing literature and industry best practices for adaptation to MLOps. For instance, Stieler and Bauer [136] proposed a Git workflow for an “active learning life cycle”, which contains four interconnected iterations of data management, model training, software development, and system operations. In the proposed workflow, a CI/CD tool hosted on a resource-rich machine with GPUs executes ML pipelines for training, testing, and integrating ML models based on the code committed by developers to a source control management system. Alternative, generic workflow engines such as Apache Airflow [^6] https://airflow.apache.org/use-cases/mlops/ can be adopted to MLOps pipelines. For instance, the FLScalize [172], an MLOps platform for federated learning, utilizes Airflow to orchestrate the interaction sequences among federated learning servers and clients as well as other monitoring workflows. Finally, there have been more ML-specific workflow orchestration tools (e.g., SageMaker Pipelines, IBM Watson Pipelines, Google Vertex AI Pipelines, Netflix Metaflow, ZenML) for defining and running MLOps pipelines.
@@ 4.7. Abstracting MLOps Complexity
State-of-the-art MLOps practices and techniques prioritize automation and optimization of MLOps lifecycle, service-oriented deployment, artifact tracking, and collaboration. Recent practices increasingly focus on domain-specific reference architectures tailored for particular contexts, systematic artifact management emphasizing provenance and lineage tracking, and frameworks and practices that support addressing compliance, fairness, and security. These developments illustrate a clear evolution toward more structured, scalable, and comprehensive MLOps implementations, capable of addressing emerging complexities in ML operationalization across diverse application domains.
@@
5. RQ3: What are the Challenges Practitioners Face when Adopting MLOps, and What Solutions have been Proposed to Address them?
@@
In RQ2 we discuss work practices, techniques and technologies that organizations need to implement to enable MLOps. However, there are challenges that hinder teams from efficiently adopting MLOps into their organization’s ML-based software system productionalization process. Thus, we identify such challenges and highlight reported solutions from the literature, and summarize them.
@@ 5.3. MLOps Platform Related Challenges
MLOps platforms are designed for end-to-end ML lifecycle management, as a means of abstracting the complexity of implementing MLOps as discussed in 4.7. This section highlights challenges and limitations related to creating and using such platforms covering platform provider and practitioner perspectives.
@@ 5.3.1. Platform customization
For companies looking to build and deploy ML models, the main goal is to generate business value from data. Thus, for most of the practitioners, the underlying MLOps platform is only instrumental and their aim is to minimize the effort on implementing and managing the platform. Cloud-based managed MLOps platforms like Sagemaker are advantageous to meet this requirement. However, using fully managed MLOps platforms means the practitioners would lose fine-grained control over the technology stack and the services supported by the platform may not meet specialized requirements of many businesses [108]. This drives practitioners to implement and manage their own platforms which add additional complexity and cost. Practitioner expectation in this scenario is easy customization of managed platforms by integrating open-source, proprietary or homegrown tools. To meet this requirement managed MLOps platform providers need to design and implement their platforms following a modular architecture, integration and customization support to enable practitioners to integrate open-source tools. As an example, AWS highlights the design of Amazon SageMaker which provides practitioners the flexibility to integrate their own workflows with open-source APIs supporting TensorFlow, PyTorch, Kubeflow, etc. [108].